i have a problem in solving the age problem
my question is
A father is three times as old as his son but after 22 years he will be only twice as old as his son find the ages of both


Answer by Alan3354(69443)

A father is three times as old as his son but after 22 years he will be only twice as old as his son find the ages of both
-------------
F = 3S
F + 22 = 2(S+22)
---
3S + 22 = 2S + 44
S = 22